OpenStudy (anonymous):

A total of 764 tickets were sold for the school play. They were either adult tickets or student tickets. There were 64 more student tickets sold than adult tickets. How many adult tickets were sold?

OpenStudy (anonymous):

if the tickets for adult were sold x (assumption) then tickets for student is x+64 so x+x+64=764 2x+64=764 2x=764-64 2x=700 x=700/2 x=350

OpenStudy (anonymous):

A garden table and a bench cost $750 combined. The cost of the garden table is two times the cost of the bench. What is the cost of the bench?

OpenStudy (anonymous):

@ayeshaafzal221 can you help me with this one

OpenStudy (anonymous):

Total = $750 750/3 = 250 250x2 = 500 The table costs $500. The bench costs $250.

OpenStudy (anonymous):

did u get it?

OpenStudy (anonymous):

yes thank you!!